# Counterfactual Sensitivity in Quantitative Spatial Models

### v2

@@ 2.2 Incorporating Parameter Uncertainty
In practice it is common to assume there is no uncertainty around estimators for the structural parameter $\theta$, both in the case where estimators are taken from other papers and in the case where estimators depend on the data. An exception is Adao, Costinot, and Donaldson (2017), which reports $\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]$ for the counterfactual change variables of interest.
@@
A leading case is where the sampling distribution is approximately normal with variance-covariance matrix $\tilde{\Sigma}\left(\tilde{\theta}\left(X_{O},N_{O}\right)\right)$ and $g_{\hat{k}}$ is smooth in $\theta$ so that we can use the delta method. [^2] This notation nests the scenario where we use an estimator from another study that used different data. In that case $\tilde{\theta}$ is independent from $\left(X_{O},N_{O}\right)$ and we would write $\tilde{\Sigma}\left(\tilde{\theta}\right)$. If $g_{\hat{k}}$ is not smooth in $\theta$ one can use the bootstrap by obtaining draws $\tilde{\theta}_{b}\sim\mathcal{N}\left(\tilde{\theta},\tilde{\Sigma}\left(\tilde{\theta}\left(X_{O},N_{O}\right)\right)\right)$ and calculating $\tilde{k}_{b}^{E}=g_{\hat{k}}\left(X_{O},N_{O};\tilde{\theta}_{b}\right)$ for $b=1,...,B$. We can then obtain a $100\left(1-\beta\right)\%$ frequentist confidence region by finding the $\beta/2$ and $21-\beta/$ 2 quantiles of $\left\{\tilde{k}_{b}^{E}\right\}$. This yields
$\tilde{k}^{E}\sim\mathcal{N}\left(\hat{k}^{E},\frac{\partial g_{\hat{k}}}{\delta\theta}|_{\theta=\tilde{\theta}\left(X_{O},N_{O}\right)}\tilde{\Sigma}\left(\tilde{\theta}\left(X_{O},N_{O}\right)\right)\frac{\partial g_{\hat{k}}}{\delta\theta}|_{\theta=\tilde{\theta}\left(X_{O},N_{O}\right)}^{{}^{\prime}}\right)\equiv\mathcal{N}\left(\hat{k}^{E},\tilde{\Sigma}\left(\tilde{k}^{E}\right)\right),$
from which we can obtain the $100\left(1-\beta\right)\%$ frequentist confidence region
$\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]=\left[\tilde{k}^{E}-z_{1-\beta/2}\sqrt{\tilde{\Sigma}\left(\tilde{k}^{E}\right)},\tilde{k}^{E}+z_{1-\beta/2}\sqrt{\tilde{\Sigma}\left(\tilde{k}^{E}\right)}\right].$
@@ 2.2.1 Running Example: Armington Model (Continued)
For the Armington model, trade flows $\left\{F_{ij}\right\}$ are the underlying data that determine the expenditure shares through $\lambda_{ij}=\frac{F_{ij}}{\sum_{\ell}F_{\ell j}}$. These trade flows are also used to estimate the trade elasticity $\varepsilon$. It follows that our object of interest is
$\hat{C}_{1}^{E}=g_{\hat{C}_{1}^{E}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\}\right)\equiv g_{\hat{C}_{1}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\};\tilde{\varepsilon}\left(\left\{F_{ij}\right\}\right)\right).$
Here, as for example outlined in Allen, Arkolakis, and Takahashi (2020), $\tilde{\varepsilon}\left(\left\{F_{ij}\right\}\right)$ is an estimator obtained from
$\log\frac{F_{ij}}{F_{jj}}=-\varepsilon\log\tilde{\tau}_{ij}+\gamma_{i}+\gamma_{j}+\phi_{ij},$ (7)
for $\tilde{\tau}_{ij}$ a proxy for the trade costs between country $i$ and $j$, $\gamma_{i}$ the country fixed effect of country $i$ and $\phi_{ij}$ an error term. [^3] An example for such a proxy is given in Allen, Arkolakis, and Takahashi (2020), which assumes that $-\varepsilon\log\tau_{ij}$ can be written as a function of their continents of origin and destination and the decile of distance between countries $i$ and $j$. From this regression, we can also learn $\tilde{\Sigma}\left(\tilde{\varepsilon}\left(\left\{F_{ij}\right\}\right)\right)$ and in turn find $\left[L_{\hat{C}_{1}^{E}}^{\beta},U_{\hat{C}_{1}^{E}}^{\beta}\right]$. It follows that our relevant $g$-functions are
$L_{\hat{C}_{1}^{E}}^{\beta}$ $=g_{L_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\}\right)\quad\mathrm{and}\quad U_{\hat{C}_{1}^{E}}^{\beta}=g_{U_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\}\right).$ (8)
3 Empirical Bayes Uncertainty Quantification
This section introduces measurement error to quantitative spatial models and discusses the differences with the existing literature on measurement error. It introduces an empirical Bayes approach to uncertainty quantification in the presence of measurement error, and outlines how to obtain credible sets for the counterfactual change variable of interest.
@@ 3.1 Introducing Measurement Error
Section 2 argued that our object of interest could be either $\hat{k}$, $\hat{k}^{E}$, $L_{\hat{k}^{E}}^{\beta}$ or $U_{\hat{k}^{E}}^{\beta}$. Furthermore, under Assumption 1, we saw that we can write the object of interest as a function solely of the observables $\left(X_{O},N_{O}\right)$. Hence, going forward, I will use the notation
$T=g_{T}\left(X_{O},N_{O}\right),$ (9)
where $g_{T}$ is a known function and, depending on the context, $T$ can equal $\hat{k}$, $\hat{k}^{E}$, $L_{\hat{k}^{E}}^{\beta}$ or $U_{\hat{k}^{E}}^{\beta}$.
The fact that $T$ only depends on the observables is convenient for answering counterfactual questions. However, the observables are economic variables which are often measured with error. For instance, Ortiz-Ospina and Beltekian (2018) and Goes (2023) highlight that there are large discrepancies between and within various data sources from trade and international economics. Motivated by this I assume that, instead of the true $X_{O}$ and $N_{O}$, we observe noisy versions $\tilde{X}_{O}$ and $\tilde{N}_{O}$, respectively, where
@@
for some independent measurement error variable $\xi\in\Xi\subseteq\mathbb{R}^{d_{\xi}}$ and measurement error function $m:\mathcal{X}_{O}\times\mathcal{N}_{O}\times\Xi\rightarrow\mathcal{X}_{O}\times\mathcal{N}_{O}$. I will assume that $m\left(\cdot\right)$ is known and will treat the distribution of $\xi$ as something to be learned.
Since the object of interest $T$ is a function of the true data, we will have uncertainty around $T$ when the data are noisily measured:
$\tilde{T}=g_{T}\left(\tilde{X}_{O},\tilde{N}_{O}\right).$
Our goal is to quantify uncertainty about $T$, and in particular to provide sets that cover $T$ with prescribed probability.
@@ 3.1.1 Relation to Measurement Error Literature
The literature on measurement error in nonlinear models is extensive, as reviewed in Hu (2015) and Schennach (2016), and the most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Hu and Schennach, 2008; Schennach, White, and Chalak, 2012; Song, Schennach, and White, 2015). However, conventional results do not apply to my setting. The key distinguishing feature of the setting in this paper is that the object of interest $T=g_{T}\left(X_{O},N_{O};\theta\right)$ directly depends on the correctly measured data, because the equality in Assumption 1 is an exact statement. In contrast, in conventional measurement error settings the object of interest is a function of the correctly measured distribution of the data, $\mathcal{P}_{O}$, rather than the actual observations, $\left(X_{O},N_{O}\right)$. This distinction is important because in my setting, even if we could perfectly estimate $\mathcal{P}_{O}$, measurement error would still be a problem. For example in the Armington model, to answer counterfactual questions we need the realized trade flows, rather than the distribution of trade flows. By virtue of that, we need to account for uncertainty about the underlying observations themselves, which is a different problem than that in the conventional measurement error literature. This reliance on the individual observations rather than the distribution is a key distinction, and it is caused by the underlying assumption that the model exactly matches the data.
3.1.2 Running Example: Armington Model (Continued)
For the Armington model, I will assume that there is measurement error in bilateral trade flows $\left\{F_{ij}\right\}$. This is plausible because in Linsi, Burgoon, and Mügge (2023) it is shown that there are so-called mirror discrepancies in bilateral trade flows between almost all countries. This means that, for instance, while the value that Germany reports it imported from France and the value that France reports it exported to Germany should be the same, in practice they are often different. Instead of the true trade flows we observe noisy trade flows $\left\{\tilde{F}_{ij}\right\}$ which then lead to noisy expenditure shares $\left\{\tilde{\lambda}_{ij}\right\}$. Since trade flows are non-negative, I assume measurement error is log-normally distributed:
$\log\tilde{F}_{ij}|\log F_{ij}\sim\mathcal{N}\left(\log F_{ij},\varsigma^{2}\right),\qquad i,j=1,...,n.$
In the case where we also consider estimation error, this means that instead of $L_{\hat{C}_{1}^{E}}^{\beta}$ and $U_{\hat{C}_{1}^{E}}^{\beta}$ we observe
$\tilde{L}_{\hat{C}_{1}^{E}}^{\beta}$ $=g_{L_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{\tilde{F}_{ij}\right\},\left\{Y_{i}\right\}\right)\quad\mathrm{and}\quad\tilde{U}_{\hat{C}_{1}^{E}}^{\beta}=g_{U_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{\tilde{F}_{ij}\right\},\left\{Y_{i}\right\}\right).$ (11)
3.2 Motivation for Empirical Bayes Approach
The previous section introduced measurement error to quantitative spatial models. We are interested in $T=g_{T}\left(X_{O},N_{O}\right)$ but instead observe $\tilde{T}=g_{T}\left(\tilde{X}_{O},\tilde{N}_{O}\right)$, which leads to the goal of generating an interval which covers $T$ with prescribed probability.
To achieve this goal, I take an empirical Bayes (EB) approach, and introduce a model for the measurement error and estimate a prior for the true underlying data. [^4] Rather than estimating the parameters of the prior for the true underlying data, which corresponds to an EB approach, one could alternatively specify priors for these parameters, which corresponds to a hierarchical Bayes approach. Using Bayes’ rule, the measurement error model and the prior can be combined to find the posterior of the true data given the noisy data. We can then feed this posterior through $T=g_{T}\left(X_{O},N_{O}\right)$ to obtain the posterior of the counterfactual change variable of interest.
A first reason for using an EB approach is that it allows researchers to incorporate economic knowledge through the prior. For example when considering measurement error in positive flows between locations one can fit a gravity model as a prior, which I will do in Section 3.4. A second reason is that, given the prior, the resulting procedure is automated and computationally appealing.
An alternative approach, which I call an "adversarial approach" following the machine learning literature, and which is developed in Appendix C, prespecifies a bound on the measurement error and generates worst-case bounds on the counterfactual. This avoids the need to fully specify a prior distribution for the measurement error, at the cost of having to prespecify a bound on the measurement error. The adversarial approach is more computationally demanding and, to be feasible in high-dimensional applications, must rely on approximations.
3.3 Obtaining Credible Sets for $T$
I now discuss how to obtain credible sets for $T=g_{T}\left(X_{O},N_{O}\right)$ when we observe $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$. That is, denoting with $\mathbb{P}_{\pi}$ the probability under the prior, for a given $\gamma\in\left[0,1\right]$ we want to find a region $\mathcal{C}_{T}^{\gamma}$ such that $\mathbb{P}_{\pi}\left[T\in\mathcal{C}_{T}^{\gamma}|\tilde{X}_{O},\tilde{N}_{O}\right]\geq 1-\gamma.$ Towards that end, given a prior on the true data and a model for the measurement error,
we can use Bayes’ rule to obtain the posterior of the true data given the noisy data
[…]
This posterior then allows us to generate draws from our posterior for the true data given the noisy data. For each of these draws, we can calculate the corresponding value for the counterfactual change variable of interest using the relationship $T=g_{T}\left(X_{O},N_{O}\right)$. This allows us to find the posterior of the true object of interest $T$ given the noisy data,
$\pi\left(T|\tilde{X}_{O},\tilde{N}_{O}\right).$
Once we have this posterior, we can obtain credible sets by for example taking an interval between two relevant quantiles. Note that these credible sets can not be interpreted as frequentist confidence intervals, and are dependent on the specified prior distribution.
Proposition 1.
Denote with $q_{\alpha}$ the $\alpha$-quantile of the posterior $T|\tilde{X}_{O},\tilde{N}_{O}$. A $100\left(1-\gamma\right)\%$ equal-tailed credible set is given by $\mathcal{C}_{T}^{\gamma}=\left[q_{\gamma/2},q_{1-\gamma/2}\right]$:
$\mathbb{P}_{\pi}\left[T\in\mathcal{C}_{T}^{\gamma}|\tilde{X}_{O},\tilde{N}_{O}\right]\geq 1-\gamma.$
Remark 1.
If we want to account for both estimation error and measurement error simultaneously, we focus on covering the $100\left(1-\beta\right)\%$ frequentist confidence region $\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]$ with prescribed probability. We then first choose $T$ equal to $L_{\hat{k}^{E}}^{\beta}$ and then to $U_{\hat{k}^{E}}^{\beta}$. Denote with $\ell_{\alpha}^{\beta}$ and $\upsilon_{\alpha}^{\beta}$ the $\alpha$-quantiles of the resulting posteriors $L_{\hat{k}^{E}}^{\beta}|\tilde{X}_{O},\tilde{N}_{O}$ and $U_{\hat{k}^{E}}^{\beta}|\tilde{X}_{O},\tilde{N}_{O}$, respectively. In that case, a $100\left(1-\beta-\gamma\right)\%$ equal-tailed credible set is given by $\mathcal{C}_{\hat{k}^{E}}^{\beta,\gamma}=\left[\ell_{\gamma/2}^{\beta},\upsilon_{1-\gamma/2}^{\beta}\right]$.
3.4 Widely Applicable Default Approach

### v3

@@ 3.1 Introducing Estimation Error
The counterfactual change variable of interest will generally depend on the structural parameter $\theta$. In practice, we do not know the structural parameter exactly and we hence have to use an estimator $\tilde{\theta}\left(X_{O},N_{O}\right)$. It is common to assume there is no uncertainty around estimators for the structural parameter $\theta$, both in the case where estimators are taken from the literature and in the case where estimators depend on the data. An exception is Adao, Costinot, and Donaldson (2017), which reports confidence sets for the counterfactual change variables of interest that account for estimation error.
I will take a Bayesian approach and assume that the posterior distribution of the true structural parameter $\theta$ given the observables $\left(X_{O},N_{O}\right)$ is approximately normal. [^2] Formally, this normality could follow from assumptions on the underlying data generating process such that a Bernstein-von Mises theorem holds (Van der Vaart, 2000). In that case the influence of the prior distribution $\pi\left(\theta\right)$ becomes negligible and the posterior distribution approximately equals a normal distribution centered at the maximum likelihood estimator. Finding appropriate sufficient conditions for a Bernstein-von Mises theorem to hold in quantitative trade and spatial models is beyond the scope of this paper. Specifically, we have
$\pi^{EE}\left(\theta|X_{O},N_{O}\right)\approx\mathcal{N}\left(\tilde{\theta}\left(X_{O},N_{O}\right),\tilde{\Sigma}\left(X_{O},N_{O}\right)\right),$
@@
where $\tilde{\Sigma}\left(X_{O},N_{O}\right)$ is a consistent estimator of the sampling variance of $\tilde{\theta}\left(X_{O},N_{O}\right)$. [^3] This notation nests the scenario where we use an estimator from another study that used different data. In that case $\theta$ is independent from $\left(X_{O},N_{O}\right)$ and we would write $\pi^{EE}\left(\theta|X_{O},N_{O}\right)\approx\mathcal{N}\left(\tilde{\theta},\tilde{\Sigma}\right)$. Furthermore, in the case where $\theta$ is known to be non-negative, one can use a log-normal distribution here.
We can generate draws from this posterior distribution of $\theta$ given $\left(X_{O},N_{O}\right)$. For each of these draws, we can calculate the corresponding value of the counterfactual change variable of interest using the relationship $\hat{k}=g_{\hat{k}}\left(X_{O},N_{O};\theta\right)$. This allows us to find the posterior distribution of $\hat{k}$ given the true data, $\pi^{EE}\left(\hat{k}|X_{O},N_{O}\right).$
@@ 3.1.1 Running Example: Armington Model (Continued)
For the Armington model, log trade flows $\left\{\log F_{ij}\right\}$ are the underlying data that determine the expenditure shares through $\lambda_{ij}=\frac{F_{ij}}{\sum_{\ell}F_{\ell j}}$, so going forward I will use $N_{O}=\left(\left\{Y_{i}\right\},\left\{\log F_{ij}\right\}\right)$. These log trade flows are also used to estimate the trade elasticity $\varepsilon$. As for example outlined in Costinot and Rodríguez-Clare (2014), $\tilde{\varepsilon}\left(\left\{\log F_{ij}\right\}\right)$ and $\tilde{\Sigma}\left(\left\{\log F_{ij}\right\}\right)$ can be obtained from the regression
$\log F_{ij}=-\varepsilon\log\tilde{\tau}_{ij}+\gamma_{i}+\gamma_{j}+\phi_{ij},$ (7)
for $\tilde{\tau}_{ij}$ a proxy for the trade costs between country $i$ and $j$, $\gamma_{i}$ the country fixed effect of country $i$ and $\phi_{ij}$ an error term that is orthogonal to $\log\tilde{\tau}_{ij}$. [^4] Note that this is a dyadic regression and the error terms are likely to be correlated. Techniques to calculate $\tilde{\Sigma}\left(\left\{\log F_{ij}\right\}\right)$ are available in for example Graham (2020). We can then find the approximate posterior distribution of our object of interest $\hat{C}_{1}$ given the true data:
$\pi^{EE}\left(\hat{C}_{1}|\left\{\log F_{ij}\right\}\right)\approx\mathcal{N}\left(\tilde{\varepsilon}\left(\left\{\log F_{ij}\right\}\right),\tilde{\Sigma}\left(\left\{\log F_{ij}\right\}\right)\right).$
@@ 3.2 Introducing Measurement Error
Under Assumption 1, our object of interest can be written as a function solely of the observables and the structural parameter, which is convenient for answering counterfactual questions. However, the observables are economic variables which are often measured with error. For instance, Ortiz-Ospina and Beltekian (2018) and Goes (2023) highlight that there are large discrepancies between and within various data sources from trade and international economics.
Motivated by this I assume that, instead of the true $X_{O}$ and $N_{O}$, we observe noisy versions $\tilde{X}_{O}$ and $\tilde{N}_{O}$, respectively, where
@@
for some independent measurement error variable $\xi\in\Xi\subseteq\mathbb{R}^{d_{\xi}}$ and measurement error function $m:\mathcal{X}_{O}\times\mathcal{N}_{O}\times\Xi\rightarrow\mathcal{X}_{O}\times\mathcal{N}_{O}$. I will assume that both $m\left(\cdot\right)$ and the distribution of $\xi$ are known, the latter possibly up to some parameters.
For uncertainty quantification about the object of interest, we will require the posterior distribution of the true data given the noisy data. Towards that end, I take an empirical Bayes (EB) approach, and introduce a model for the measurement error and estimate a prior distribution for the true underlying data. [^5] Rather than estimating the parameters of the prior distribution for the true underlying data, which corresponds to an empirical Bayes approach, one could alternatively specify prior distributions for these parameters, which corresponds to a hierarchical Bayes approach. Given such a prior distribution and a measurement error model,
we can use Bayes’ rule to find the posterior distribution of the true data given the noisy data,
[…]
This posterior distribution then allows us to generate draws from our posterior distribution for the true data given the noisy data.
The EB approach allows researchers to incorporate economic knowledge through the prior. For example when considering measurement error in positive flows between locations one can fit a gravity model as a prior, which I will do in Section 3.4. Furthermore, given the prior, the resulting procedure is automated and computationally appealing.
3.2.1 Running Example: Armington Model (Continued)
For the Armington model, I will assume that there is measurement error in log bilateral trade flows $\left\{\log F_{ij}\right\}$. This is plausible because in Linsi, Burgoon, and Mügge (2023) it is shown that there are so-called mirror discrepancies in bilateral trade flows between almost all countries. This means that, for instance, while the value that Germany reports it imported from France and the value that France reports it exported to Germany should be the same, in practice they are often different. Instead of the true log trade flows we observe noisy trade flows $\left\{\log\tilde{F}_{ij}\right\}$ which then lead to noisy expenditure shares $\left\{\tilde{\lambda}_{ij}\right\}$. If we specify a prior $\pi\left(\left\{\log F_{ij}\right\}\right)$ and a measurement error model $\pi\left(\left\{\log\tilde{F}_{ij}\right\}|\left\{\log F_{ij}\right\}\right)$, we can use Bayes’ rule to find the posterior $\pi^{ME}\left(\left\{\log F_{ij}\right\}|\left\{\log\tilde{F}_{ij}\right\}\right)$.
@@ 3.2.2 Relation to Measurement Error Literature
The literature on measurement error in nonlinear models is extensive, as reviewed in Hu (2015) and Schennach (2016), and the most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Hu and Schennach, 2008; Schennach, White, and Chalak, 2012; Song, Schennach, and White, 2015). However, these results do not apply to my setting. The key distinguishing feature of the setting in this paper is that the object of interest $\hat{k}$ directly depends on the correctly measured data, because the equality in Assumption 1 is an exact statement. In contrast, in conventional measurement error settings the object of interest is a function of the correctly measured distribution of the data, $\mathcal{P}_{O}$, rather than the actual observations, $\left(X_{O},N_{O}\right)$. This distinction is important because in my setting, even if we could perfectly estimate $\mathcal{P}_{O}$, measurement error would still be a problem. For example in the Armington model, to answer counterfactual questions we need the realized trade flows, rather than the distribution of trade flows. By virtue of that, we need to account for uncertainty about the underlying observations themselves, which is a different problem than that in the conventional measurement error literature. This reliance on the individual observations rather than the distribution is a key distinction, and it stems from the underlying assumption that the model exactly matches the data.
3.3 Quantifying Uncertainty about $\hat{k}$
The object of interest is a function of the true data and the structural parameter. From the discussion in the previous sections, it follows that we must consider estimation error, the direct effect of mismeasurement, and the indirect effect of mismeasurement through the estimation procedure. Our goal is to quantify uncertainty about $\hat{k}$ when we observe $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ by accounting for these various sources of uncertainty.
Recall that we have obtained two different posteriors. The first one is the posterior distribution of $\hat{k}$ given the true data, $\pi^{EE}\left(\hat{k}|X_{O},N_{O}\right)$, which incorporates estimation error. The second one is the posterior of the true data given the noisy data, $\pi^{ME}\left(X_{O},N_{O}|\tilde{X}_{O},\tilde{N}_{O}\right)$, which incorporates measurement error. We can combine these two posteriors in two different ways to quantify uncertainty about $\hat{k}$. Here, I will use $\mathbb{E}_{\pi}$ and $Pr_{\pi}$ to denote the expectation and probability under a posterior $\pi$, respectively.
The first approach aims to find an interval $\mathcal{C}^{1}$ to which, in posterior expectation over $\left(X_{O},N_{O}\right)$, the posterior $\pi^{EE}\left(\hat{k}|X_{O},N_{O}\right)$ assigns probability $1-\alpha$:
$\mathbb{E}_{\pi^{ME}}\left[Pr_{\pi^{EE}}\left\{\hat{k}\in\mathcal{C}^{1}|X_{O},N_{O}\right\}|\tilde{X}_{O},\tilde{N}_{O}\right]\geq 1-\alpha.$
In practice, given $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ one would generate draws from $\pi^{ME}\left(X_{O},N_{O}|\tilde{X}_{O},\tilde{N}_{O}\right)$, and for each of these draws obtain a corresponding draw from $\pi^{EE}\left(\hat{k}|X_{O},N_{O}\right)$. Then, one would report the $\alpha/2$ and $21-\alpha/$ 2 quantiles of this second set of draws.
The second approach is more conservative. Suppose we again obtain draws from the posterior $\pi^{ME}\left(X_{O},N_{O}|\tilde{X}_{O},\tilde{N}_{O}\right)$ and for each draw use $\pi^{EE}\left(\hat{k}|X_{O},N_{O}\right)$ to compute an interval that covers $\hat{k}$ with probability $1-\alpha$. The second approach then aims to find an interval $\mathcal{C}^{2}$ that covers these $100\left(1-\alpha\right)\%$intervals with probability $1-\alpha$:
$Pr_{\pi^{ME}}\left\{Pr_{\pi^{EE}}\left\{\hat{k}\in\mathcal{C}^{2}|X_{O},N_{O}\right\}\geq 1-\alpha|\tilde{X}_{O},\tilde{N}_{O}\right\}\geq 1-\alpha.$
In practice, one can generate many $100\left(1-\alpha\right)\%$intervals around $\hat{k}$ and report the $\alpha/2$ quantile of the set of lower bounds and the $21-\alpha/$ 2 quantile of the set of upper bounds.
Going forward, I will focus on the less conservative interval $\mathcal{C}^{1}$, because it will turn out that in applications the interval $\mathcal{C}^{2}$ will yield extremely wide intervals for many cases. In Appendix C.3 I compute the interval $\mathcal{C}^{2}$ for one of my applications. [^6] Note that one could in principle use a single prior $\pi$ on the underlying data generating process to handle both estimation error and measurement error. I instead combine two simple priors to separately handle estimation error and measurement error, since this leads to highly tractable procedures, albeit at the cost of complicating the Bayesian interpretation of resulting intervals.
3.4 Widely Applicable Default Approach for Uncertainty Quantification
In this section I provide a widely applicable default approach for quantifying uncertainty about a counterfactual change variable of interest. This default approach can be applied out-of-the-box to many quantitative trade and spatial models, but can also easily be adapted to other settings. It recommends default choices for the prior distribution and measurement error model, and discusses how to calibrate both based on observed data.